OpenStudy (anonymous):

FOR THE LOVE OF PETE WENTZ CAN SOMEBODY PLEASE HELP ME SOLVE THIS did the title get your attention? 4g^2 - 32g = 0

OpenStudy (anonymous):

you sure have been typing for a long time.....

OpenStudy (owlcoffee):

In these types of equations, we can apply what is called "common factor" or as you might, divide and multiply the whole expression by one convenient rational number that includes "g". \[4g^2-32g=0 \iff \frac{ 4g }{ 4g }(4g^2-32g)=0 \] This is a valid transformation for two reasons, that \(1(4g^2-32g)=(4g^2-32g)\) and because \(\frac{ 4g }{ 4g } =1\), which means that we have not altered the equation, just rewritten some values to make the work more simple, so, by distributive property we obtain: \[4g(\frac{ 4g^2 }{ 4g }-\frac{ 32g }{ 4g })=0 \iff 4g(g-8)=0\] From the hankelian property, we can fragment this las equality into two much more simple expressions: \[4g=0 \] \[g-8=0 \] So, solving these two equations will solve the original problem equation.

OpenStudy (anonymous):

the first one isn't really a valid answer thought is it? because 0 can't be divided. so the only real answer to this would be g=8 right?

OpenStudy (anonymous):

(btw that was amazing)

OpenStudy (owlcoffee):

Oh, zero can be divided, think of it this way, if you have no candies and you want to share it with 7 friends, how many do each get? The answer is no one gets any candy, so, any number that divides zero will yield a result of zero. \[\frac{ 0 }{ k }=0 \] Problems begin when zero is the one to divide a number, implying, that you have candies, but you have to share them with no one , how many candies does each person get? Nobody really knows, but in mathematics we don't agree on a result, we prove it and no one has really figures what any number divided by zero is equal to. \[\frac{ k }{ 0 }=undef.\]